OpenStudy (anonymous):

The measures of the angles of a triangle are in the extended ratio 2:2:8. What is the measure of the smallest angle?

OpenStudy (alexandervonhumboldt2):

o use this: 2+2+8=12 there are 12 parts

OpenStudy (alexandervonhumboldt2):

the goal is to find the first part. the sum of all interior angles of a triangle is 180

OpenStudy (alexandervonhumboldt2):

so divide 180/12=15

OpenStudy (alexandervonhumboldt2):

now multiply 15*2=30 so 30 degrees is the smallest angle

OpenStudy (anonymous):

Thank you so much :)

OpenStudy (alexandervonhumboldt2):

cn you plz medal and fan me and if you like write a testimonial?